OpenStudy (anonymous):

Factor k2 + 13k + 12.

OpenStudy (amistre64):

k^2 + (b+c) k + (b*c) is what we are looking at; and we want to find the value of b and c

OpenStudy (amistre64):

what 2 numbers multiply to get 12 and add to get 13? bc = 12 b+c = 13

OpenStudy (amistre64):

the rest of it boils down to trial and error 1*12 = 12 ; 1+12 = 13 .... and that dint take to long lol

OpenStudy (amistre64):

(k+b) (k+c) is how it factors out (k+12) (k+1)